OpenStudy (anonymous):

What is the value of b2 - 4ac for the following equation x 2 + 5x + 4 = 0 9 16 25

OpenStudy (jonnyvonny):

a = x^2 b = 5x c = 4 5^2 - 4(1)(4) = 25-16=?

OpenStudy (anonymous):

9?

OpenStudy (jonnyvonny):

Yes!